What minimum heat is needed to bring 100 g of water at 50 ∘C to the boiling point and completely boil it away? The specific heat of water is 4190 J/(kg⋅K) and its heat of vaporization is 22.6×105 J/kg. Express your answer to two significant figures and include the appropriate units.
Answer #1

mass of water m=100g=0.1 kg.

Sp. Heat s=4190J/(kg. K)

To rise the temperature form 50C to 100Cis t=100- 50=50C

The required heat to raise the temp from 50C to 100C Q=mst= 0.1 × 4190 × 50 = 20950 J

Heat requires to vaporise the same water = 0.1× 22.6 × 105 = 22.5 × 104 = 225000J

Total heat rsquired 225000 + 20950=245950 J (Ans.)